A solution is prepared by combining 38.614 mL of 2.10 M aqueous weak acid and 38.614 mL of 1.05 M aqueous NaOH. What is the pH of the resulting solution? The Ka of the weak acid is 8.844 x 10-8. Enter your answer with at least two decimal places.

Please show your work for how you got your answer.
Answer #1

Answer:

Volume of acid = 38.614 mL = 38.614 × 10-3 L

Concentration of acid = 2.10M

Ka = 8.844 × 10-8

[H+]= (Ka×C)1/2

[H+]= (8.844 × 10-8×2.10)1/2

[H+]= 4.309 × 10-4 M

No. of equivalents of H+ = [H+]× V

No. of equivalents of H+ = 4.309 × 10-4 × 38.614 × 10-3

No. of equivalents of H+ =1.664 × 10-5

Volume of NaOH = 38.614 mL = 38.614 × 10-3 L

Concentration of acid = 1.05 M

No. of equivalents of OH- = C× V

No. of equivalents of OH- = 1.05 × 38.614 × 10-3

No. of equivalents of OH- =4.045 × 10-2

Out of which,1.664 × 10-5 OH- ions will be neutralised by H+ ions.

OH- ions remaining in the solution = 4.045 × 10-2 - 1.664 × 10-5

OH- ions remaining in the solution = 0.04053121

POH = -log[OH-]

POH = -log[0.04053121]

POH = 1.3922

PH = 14-POH

PH = 14-1.3922

PH = 12.60